# AudioMarkBench: Benchmarking Robustness of Audio Watermarking

### v1

@@ 1 Introduction
Recent advancements in text-to-speech (TTS) generative models enable generating highly realistic synthetic audios that are indistinguishable from real human voices. However, this capability raises significant concerns, such as malicious impersonation, dissemination of false information, or copyright infringement. For example, a scammer used synthetic audios to impersonate President Biden in illegal robocalls during a New Hampshire primary election, and thus faces a \$6 million fine and felony charges [5].
Audio watermarking [12, 3, 10] offers a promising approach to mitigate concerns about synthetic audio authenticity. It embeds an imperceptible watermark into a synthetic audio using a watermark encoder, outputting a watermarked audio. During detection, a watermark decoder extracts a watermark from a given audio input. By comparing the extracted watermark with the ground-truth watermark, one can determine the authenticity of the given audio.
Existing audio watermarking methods perform well when there are no perturbations added to watermarked audios [12, 3, 10]. However, real-world audios often undergo various perturbations. Common perturbations include compression using standards like MP3 or Opus [13] to reduce internet transmission costs. Additionally, attackers may craft adversarial perturbations designed to deceive watermarking methods. However, the robustness of audio watermarking against these perturbations remains under-explored and lacks systematic benchmarking.
@@
- Datasets: Other than the standard LibriSpeech dataset [11], we construct a new dataset AudioMarkData that meticulously sub-samples 20,000 audio samples from the Common Voice dataset [2], striving to ensure balanced representation of biological sexes, languages, and age groups. Moreover, our datasets provide not only watermarked/unwatermarked audios but also perturbed audios under various perturbations, making it easier for future research to assess the the effectiveness of new watermark-removal/forgery perturbations.
- Systematic benchmarking: We present the first systematic benchmark evaluating the robustness of three state-of-the-art audio watermarking methods against 15 different watermark-removal/forgery perturbations across two datasets. Twelve of these perturbations, termed “no-box” perturbations, require no access to the watermarking method. These perturbations include common audio edits like codec [6, 14, 13] and audio filter, and noise addition such as white noise or background noise. Additionally, We adapt two adversarial example methods [4, 1] in the black-box setting (i.e., access to watermark detector API only) and one adversarial example method [9] in the white-box setting (i.e., full access to watermarking model parameters) from image classifiers to audio watermarking.
- Findings: We make intriguing findings in our benchmark study. First, we confirm that all studied audio watermarking methods can distinguish watermarked/AI-generated audios from unwatermarked/non-AI-generated audios precisely when no perturbations are added. Second, existing audio watermarking methods can be vulnerable to watermark removal including certain no-box perturbations (e.g., EnCodeC [6]), black-box perturbations with sufficient quota for API queries, and white-box perturbations. Third, current audio watermarking techniques are effective at resisting no-box and black-box watermark forgery, but vulnerable to white-box forgery. Fourth, existing audio watermarking methods have robustness gaps among biological sex groups (female vs male) and language groups under certain perturbations, flagging potential fairness issues. However, we do not observe consistently significant robustness gaps across age groups.
@@ 2 Audio Watermarking Methods
An audio watermarking method consists of four key components: a watermark $w$, encoder Enc, decoder Dec, and detector Det. The watermark $w\in\{0,1\}^{n}$ is typically an $n$-bit bitstring, such as an 16-bit bitstring $1110110110010110$. Given any audio waveform $s\in\mathbb{R}^{T}$ and a watermark $w$, the encoder Enc outputs a watermarked audio waveform $s_{w}=\texttt{Enc}(w,s)\in\mathbb{R}^{T}$, where $T$ denotes the number of time samples in the waveform. For any audio waveform $s$, whether watermarked or unwatermarked, the decoder Dec can extract a bitstring watermark $\texttt{Dec}(s)$. When the audio waveform $s$ is watermarked with $w$, the extracted watermark $\texttt{Dec}(s)$ should be similar to $w$. The detector Det then uses the decoded watermark $\texttt{Dec}(s)$, together with some additional information, to determine if the given audio waveform $s$ contains a watermark. In particular, $\texttt{Det}(s)=1$ ($\texttt{Det}(s)=0$) means that $s$ is detected as watermarked (unwatermarked), respectively. In this study, we examine three state-of-the-art, open-source audio watermarking techniques: AudioSeal/AudioSeal-B [12], Timbre [10], and WavMark [3]. We utilize the publicly available code and models of these methods for our experimental analysis.
@@ 3 Watermark-removal and Watermark-forgery Perturbations
where $Q$ is an audio quality metric. The quality constraint ensures the audio quality to remain high after adding the perturbation. The audio quality metric $Q$ can be ViSQOL [8] or SNR.
@@
Black-box perturbations:  In black-box setting, perturbations are created by interacting with the watermarking detector Det as an oracle. Specifically, the attacker can choose audios to submit to the detector and observe the detection result without any knowledge of how the detector operates internally. We extend existing methods for finding black-box adversarial examples [4, 1] against image classifiers to audio watermarking detectors. In particular, we apply them in waveform and/or spectrogram domains. Next, we briefly describe how we extend them, and Appendix A.3 shows more technical details.
- HopSkipJumpAttack (HSJA) [4]: Given an audio $s$ and access to a watermarking detector Det’s output, HopSkipJumpAttack iteratively approximates Det’s decision boundary to find a minimal watermark-removal/watermark-forgery perturbation $\delta$. We implement this attack in both waveform and spectrogram domains. In the waveform domain, perturbations are optimized in a 1-D vector space, while in the spectrogram domain, both phase and amplitude (2-D vectors) are optimized. We conduct 10,000 iterations in each domain, initializing perturbations with Gaussian noise.
- Square attack [1]: Given an audio $s$ and access to a watermarking decoder Dec’s output, Square attack iteratively finds the watermark-removal/watermark-forgery perturbation $\delta$ by strategically decreasing either the bitwise accuracy of Dec’s output or the global detection probability (for AudioSeal). We extend Square attack from image domain to the spectrogram domain by treating a spectrogram as an image. Note that Square attack is only applicable to the spectrogram domain (not the waveform domain) since its input is a 2-D image/spectrogram. We perform Square attack under a $\ell_{\infty}$-norm perturbation constraint for 10,000 iterations.
@@
$L_{ce}=-\sum_{i=1}^{n}w_{i}\log(\texttt{Dec}(s+\delta)_{i})+(1-w_{i})\log(1-\texttt{Dec}(s+\delta)_{i}),$
where $w_{i}$ (or $\texttt{Dec}(s+\delta)_{i}$) is the $i^{th}$ bit of $w$ (or $\texttt{Dec}(s+\delta)$). For AudioSeal, we adopt the ReLU activation applied between the global detection probability $P_{s}$ and $\tau$: $L_{Re}=\max(0,P_{s}-\tau)$. We use these loss functions to approximate the objective functions in Equation 1 and 2. Appendix A.4 shows more details on how we solve the optimization problems to find white-box perturbations.
@@ 4 Datasets
Unwatermarked audio samples:  Our AudioMarkBench includes two datasets of unwatermarked audio samples, i.e., AudioMarkData and LibriSpeech [11]. AudioMarkData is a dataset we build from the Common Voice dataset [2]. Each audio sample in AudioMarkData is associated with three attributes, which are: language (25 languages), biological sex (male, female) and age (teens, twenties, thirties, fourties). We use these attributes to benchmark whether watermarking methods have different performance/robustness for audio samples with different attributes. For every attribute group (language, biological sex, age), AudioMarkData samples 100 audio samples in 5 seconds with sampling rate at 16kHz from Common Voice, resulting in 20,000 audio samples in total. Table 1 summarizes the attributes of AudioMarkData. The LibriSpeech dataset contains over 1,000 hours of read English speech derived from audiobooks in the public domain. We sampled 20,000 audio samples with a maximum length of 5 seconds at the default 16kHz sampling rate. Note that audio samples in LibriSpeech do not have attributes.
@@ 5 Benchmark Results
In the following section, we present our primary benchmark results and findings. We conduct our experiments on 18 NVIDIA-RTX-6000 GPUs, each with 24 GB memory. All experiments requires about 430 GPU-hours to excute.
@@ 5.1 Evaluation Metrics
We evaluate the quality of perturbed audios using standard metrics including SNR and ViSQOL [8]. Signal-to-Noise Ratio (SNR) evaluates quality of a perturbed audio by comparing its level of noise with the corresponding clean audio (called reference audio), where the reference audio is watermarked (or unwatermarked) in watermark removal (or forgery). Higher SNRs indicate clearer and higher-quality perturbed audios. ViSQOL, ranging from 1 to 5, evaluates audio quality by simulating human perception of audios, where a higher score indicates the perturbed audio better preserves quality of the reference audio. A ViSQOL score no smaller than 3 generally reflects good audio quality. We mainly rely on ViSQOL for measuring audio quality because it is more reliable than SNR [8].
@@ 5.3 Robustness against No-box Perturbations
Overall results: We have several key observations. First, state-of-the-art audio watermarks are robust against several common no-box watermark-removal perturbations such as time stretch, low-pass, high-pass, and echo. Specifically, while preserving the quality of watermarked audio samples well (i.e., ViSQOL no smaller than 3), those perturbations have small impact on FNRs. This is because these audio watermarking methods use adversarial training [7], which considers various common no-box perturbations, to train the encoders and decoders. Second, current audio watermarking methods are not robust against no-box removal perturbations that are unseen during adversarial training. For instance, when ViSQOL is no smaller than 3, EnCodeC, SoundStream, and Opus achieve very high FNRs, indicating that those perturbations can remove watermarks from watermarked audios while preserving the audio quality. Third, current audio watermarking methods have good robustness against watermark-forgery perturbations. In particular, FPRs of all these watermarking methods are almost always close to 0, except for quantization. Specifically, when bit levels are smaller than 32, quantization perturbation achieves a FPR larger than 0.2, but the audio quality is also compromised. This is because forging a watermark is harder and may require knowledge of the watermarking model. No-box perturbations do not have such information and therefore cannot forge a watermark. As we will show in the next subsection, forging a watermark remains difficult even in the black-box setting.
@@ 5.4 Robustness against Black-box Perturbations
5.5 Robustness against White-box Perturbations
Figure 7: ViSQOL vs. SNR of white-box perturbations.
Figure 8 shows the detection results under white-box perturbations, where the perturbations are constrained by SNR. We evaluate SNRs from 20 to 60, which correspond to ViSQOL scores from above 3 to 5 (see Figure 7). In other words, our white-box perturbations preserve the audio quality. Our key observation is that existing audio watermarking methods are not robust to white-box watermark-removal and watermark-forgery perturbations. For instance, FNRs reach 1 for all watermarking methods when the SNR of the perturbations is 20 (i.e., ViSQOL of 3.2 and 3.9 on the two datasets). Moreover, all watermarking methods have high FPRs under white-box perturbations that preserve audio quality.

### v2

@@ 1 Introduction
Recent advancements in text-to-speech (TTS) generative models enable generating highly realistic synthetic audios that are indistinguishable from real human voices. However, this capability raises significant concerns, such as malicious impersonation, dissemination of false information, or copyright infringement. For example, a scammer used synthetic audios to impersonate President Biden in illegal robocalls during a New Hampshire primary election, and thus faces a \$6 million fine and felony charges Coldewey (2024).
Audio watermarking Roman et al. (2024); Chen et al. (2023); Liu et al. (2024) offers a promising approach to mitigate concerns about synthetic audio authenticity. It embeds an imperceptible watermark into a synthetic audio using a watermark encoder, outputting a watermarked audio. During detection, a watermark decoder extracts a watermark from a given audio input. By comparing the extracted watermark with the ground-truth watermark, one can determine the authenticity of the given audio.
Existing audio watermarking methods perform well when there are no perturbations added to watermarked audios Roman et al. (2024); Chen et al. (2023); Liu et al. (2024). However, real-world audios often undergo various perturbations. Common perturbations include compression using standards like MP3 or Opus Valin et al. (2012) to reduce internet transmission costs. Additionally, attackers may craft adversarial perturbations designed to deceive watermarking methods. However, the robustness of audio watermarking against these perturbations remains under-explored and lacks systematic benchmarking.
@@
- Datasets: Other than the standard LibriSpeech dataset Panayotov et al. (2015), we construct a new dataset AudioMarkData that meticulously sub-samples 20,000 audio samples from the Common Voice dataset Ardila et al. (2020), striving to ensure balanced representation of biological sexes, languages, and age groups. Moreover, our datasets provide not only watermarked/unwatermarked audios but also perturbed audios under various perturbations, making it easier for future research to assess the the effectiveness of new watermark-removal/forgery perturbations.
- Systematic benchmarking: We present the first systematic benchmark evaluating the robustness of three state-of-the-art audio watermarking methods against 15 different watermark-removal/forgery perturbations across two datasets. Twelve of these perturbations, termed “no-box” perturbations, require no access to the watermarking method. These perturbations include common audio edits like codec Défossez et al. (2022); Zeghidour et al. (2021); Valin et al. (2012) and audio filter, and noise addition such as white noise or background noise. Additionally, We adapt two adversarial example methods Chen et al. (2020); Andriushchenko et al. (2020) in the black-box setting (i.e., access to watermark detector API only) and one adversarial example method Jiang et al. (2023) in the white-box setting (i.e., full access to watermarking model parameters) from image classifiers to audio watermarking.
- Findings: We make intriguing findings in our benchmark study. First, we confirm that all studied audio watermarking methods can distinguish watermarked/AI-generated audios from unwatermarked/non-AI-generated audios precisely when no perturbations are added. Second, existing audio watermarking methods can be vulnerable to watermark removal including certain no-box perturbations (e.g., EnCodeC Défossez et al. (2022)), black-box perturbations with sufficient quota for API queries, and white-box perturbations. Third, current audio watermarking techniques are effective at resisting no-box and black-box watermark forgery, but vulnerable to white-box forgery. Fourth, existing audio watermarking methods have robustness gaps among biological sex groups (female vs male) and language groups under certain perturbations, flagging potential fairness issues. However, we do not observe consistently significant robustness gaps across age groups.
@@ 2 Audio Watermarking Methods
An audio watermarking method consists of four key components: a watermark $w$, encoder Enc, decoder Dec, and detector Det. The watermark $w\in\{0,1\}^{n}$ is typically an $n$-bit bitstring, such as an 16-bit bitstring $1110110110010110$. Given any audio waveform $s\in\mathbb{R}^{T}$ and a watermark $w$, the encoder Enc outputs a watermarked audio waveform $s_{w}=\texttt{Enc}(w,s)\in\mathbb{R}^{T}$, where $T$ denotes the number of time samples in the waveform. For any audio waveform $s$, whether watermarked or unwatermarked, the decoder Dec can extract a bitstring watermark $\texttt{Dec}(s)$. When the audio waveform $s$ is watermarked with $w$, the extracted watermark $\texttt{Dec}(s)$ should be similar to $w$. The detector Det then uses the decoded watermark $\texttt{Dec}(s)$, together with some additional information, to determine if the given audio waveform $s$ contains a watermark. In particular, $\texttt{Det}(s)=1$ ($\texttt{Det}(s)=0$) means that $s$ is detected as watermarked (unwatermarked), respectively. In this study, we examine three state-of-the-art, open-source audio watermarking techniques: AudioSeal/AudioSeal-B Roman et al. (2024), Timbre Liu et al. (2024), and WavMark Chen et al. (2023). We utilize the publicly available code and models of these methods for our experimental analysis.
@@ 3 Watermark-removal and Watermark-forgery Perturbations
where $Q$ is an audio quality metric. The quality constraint ensures the audio quality to remain high after adding the perturbation. The audio quality metric $Q$ can be ViSQOL Hines et al. (2015) or SNR.
@@
Black-box perturbations:  In black-box setting, perturbations are created by interacting with the watermarking detector Det as an oracle. Specifically, the attacker can choose audios to submit to the detector and observe the detection result without any knowledge of how the detector operates internally. We extend existing methods for finding black-box adversarial examples Chen et al. (2020); Andriushchenko et al. (2020) against image classifiers to audio watermarking detectors. In particular, we apply them in waveform and/or spectrogram domains. Next, we briefly describe how we extend them, and Appendix A.3 shows more technical details.
- HopSkipJumpAttack (HSJA) Chen et al. (2020): Given an audio $s$ and access to a watermarking detector Det’s output, HopSkipJumpAttack iteratively approximates Det’s decision boundary to find a minimal watermark-removal/watermark-forgery perturbation $\delta$. We implement this attack in both waveform and spectrogram domains. In the waveform domain, perturbations are optimized in a 1-D vector space, while in the spectrogram domain, both phase and amplitude (2-D vectors) are optimized. We conduct 10,000 iterations in each domain, initializing perturbations with Gaussian noise.
- Square attack Andriushchenko et al. (2020): Given an audio $s$ and access to a watermarking decoder Dec’s output, Square attack iteratively finds the watermark-removal/watermark-forgery perturbation $\delta$ by strategically decreasing either the bitwise accuracy of Dec’s output or the global detection probability (for AudioSeal). We extend Square attack from image domain to the spectrogram domain by treating a spectrogram as an image. Note that Square attack is only applicable to the spectrogram domain (not the waveform domain) since its input is a 2-D image/spectrogram. We perform Square attack under a $\ell_{\infty}$-norm perturbation constraint for 10,000 iterations.
@@
$L_{ce}=-\sum_{i=1}^{n}w_{i}\log(\texttt{Dec}(s+\delta)_{i})+(1-w_{i})\log(1-\texttt{Dec}(s+\delta)_{i})$
, where $w_{i}$ (or $\texttt{Dec}(s+\delta)_{i}$) is the $i^{th}$ bit of $w$ (or $\texttt{Dec}(s+\delta)$). For AudioSeal, we adopt the ReLU activation applied between the global detection probability $P_{s}$ and $\tau$: $L_{Re}=\max(0,P_{s}-\tau)$. We use these loss functions to approximate the objective functions in Equation 1 and 2. Appendix A.4 shows more details on how we solve the optimization problems to find white-box perturbations.
@@ 4 Datasets
Unwatermarked audio samples:  Our AudioMarkBench includes two datasets of unwatermarked audio samples, i.e., AudioMarkData and LibriSpeech Panayotov et al. (2015). AudioMarkData is a dataset we build from the Common Voice dataset Ardila et al. (2020). Each audio sample in AudioMarkData is associated with three attributes, which are: language (25 languages), biological sex (male, female) and age (teens, twenties, thirties, fourties). We use these attributes to benchmark whether watermarking methods have different performance/robustness for audio samples with different attributes. For every attribute group (language, biological sex, age), AudioMarkData samples 100 audio samples in 5 seconds with sampling rate at 16kHz from Common Voice, resulting in 20,000 audio samples in total. Table 1 summarizes the attributes of AudioMarkData. The LibriSpeech dataset contains over 1,000 hours of read English speech derived from audiobooks in the public domain. We sampled 20,000 audio samples with a maximum length of 5 seconds at the default 16kHz sampling rate. Note that audio samples in LibriSpeech do not have attributes.
@@ 5 Benchmark Results
In the following section, we present our primary benchmark results and findings. We conduct our experiments on 18 NVIDIA-RTX-6000 GPUs, each with 24 GB memory. The complete set of experiments requires about 430 GPU-hours to execute.
@@ 5.1 Evaluation Metrics
We evaluate the quality of perturbed audios using standard metrics including SNR and ViSQOL Hines et al. (2015). Signal-to-Noise Ratio (SNR) evaluates quality of a perturbed audio by comparing its level of noise with the corresponding clean audio (called reference audio), where the reference audio is watermarked (or unwatermarked) in watermark removal (or forgery). Higher SNRs indicate clearer and higher-quality perturbed audios. ViSQOL, ranging from 1 to 5, evaluates audio quality by simulating human perception of audios, where a higher score indicates the perturbed audio better preserves quality of the reference audio. A ViSQOL score no smaller than 3 generally reflects good audio quality. We mainly rely on ViSQOL for measuring audio quality because it is more reliable than SNR Hines et al. (2015).
@@ 5.3 Robustness against No-box Perturbations
Overall results: We have several key observations. First, state-of-the-art audio watermarks are robust against several common no-box watermark-removal perturbations such as time stretch, low-pass, high-pass, and echo. Specifically, while preserving the quality of watermarked audio samples well (i.e., ViSQOL no smaller than 3), those perturbations have small impact on FNRs. This is because these audio watermarking methods use adversarial training Goodfellow et al. (2014), which considers various common no-box perturbations, to train the encoders and decoders. Second, current audio watermarking methods are not robust against no-box removal perturbations that are unseen during adversarial training. For instance, when ViSQOL is no smaller than 3, EnCodeC, SoundStream, and Opus achieve very high FNRs, indicating that those perturbations can remove watermarks from watermarked audios while preserving the audio quality. Third, current audio watermarking methods have good robustness against watermark-forgery perturbations. In particular, FPRs of all these watermarking methods are almost always close to 0, except for quantization. Specifically, when bit levels are smaller than 32, quantization perturbation achieves a FPR larger than 0.2, but the audio quality is also compromised. This is because forging a watermark is harder and may require knowledge of the watermarking model. No-box perturbations do not have such information and therefore cannot forge a watermark. As we will show in the next subsection, forging a watermark remains difficult even in the black-box setting.
@@ 5.4 Robustness against Black-box Perturbations
5.5 Robustness against White-box Perturbations
Figure 8: ViSQOL vs. SNR of white-box perturbations.
Figure 7 shows the detection results under white-box perturbations, where the perturbations are constrained by SNR. We evaluate SNRs from 20 to 60, which correspond to ViSQOL scores from above 3 to 5 (see Figure 8). In other words, our white-box perturbations preserve the audio quality. Our key observation is that existing audio watermarking methods are not robust to white-box watermark-removal and watermark-forgery perturbations. For instance, FNRs reach 1 for all watermarking methods when the SNR of the perturbations is 20 (i.e., ViSQOL of 3.2 and 3.9 on the two datasets). Moreover, all watermarking methods have high FPRs under white-box perturbations that preserve audio quality. We also evaluate iterative Fast Gradient Sign Method (I-FGSM)  Kurakin et al. (2017) in Appendix A.11 and get similar conclusion.
@@ 7 Conclusion
Acknowledgments
We thank the anonymous reviewers for their constructive comments. This work was supported by NSF under Grant No. 2414406.
@@ A.9 Languages Differences against Watermark-removal Perturbations
A.10 Results on FMA Music Dataset
We conducted additional experiments using the FMA music dataset Defferrard et al. [2016]. Table 5 shows the results without any perturbation; Table 6 shows the results under several no-box perturbations; Table 7 shows the results under black-box attacks; and Table 8 and Table 9 show the results under white-box removal and forgery attacks.
Table 5: Detection results under no perturbations on FMA.
(a) AudioSeal
Tau 0.01 0.02 0.05 0.1 0.15
FPR 0.210 0.130 0.060 0.030 0.020
FNR 0.000 0.000 0.000 0.000 0.000
(b) AudioSeal-B
Tau 0.625 0.6875 0.75 0.8125 0.875
FPR 0.080 0.010 0.010 0.000 0.000
FNR 0.000 0.000 0.010 0.040 0.110
(c) Timbre
Tau 0.625 0.6875 0.75 0.8125 0.875
FPR 0.100 0.000 0.000 0.000 0.000
FNR 0.000 0.000 0.000 0.010 0.030
(d) WavMark
Tau 0.0 0.0625 0.125 0.1875 0.25
FPR 0.000 0.000 0.000 0.000 0.000
FNR 0.000 0.000 0.000 0.000 0.000
Table 6: Detection results (FNR/FPR) under background noise and time stretch on FMA.
(a) Background noise
SNR AudioSeal AudioSeal-B Timbre WavMark
5 .03/.01 .44/.00 .52/.00 .96/.00
10 .00/.02 .26/.00 .25/.00 .68/.00
20 .00/.01 .04/.00 .04/.00 .07/.00
30 .00/.01 .02/.01 .02/.00 .00/.00
40 .00/.01 .01/.00 .01/.00 .00/.00
(b) Time stretch
Stretch AudioSeal AudioSeal-B Timbre WavMark
0.7 .60/.01 .24/.01 .08/.00 .41/.00
0.9 .45/.02 .19/.00 .04/.00 .20/.00
1.1 .25/.02 .10/.01 .04/.00 .24/.00
1.3 .57/.01 .17/.01 .12/.00 .57/.00
1.5 .73/.02 .31/.02 .13/.00 .84/.00
Table 7: Results under black-box Square attack on FMA.
(a) AudioSeal
$\ell_{\infty}$ Bound FNR SNR ViSQOL
0.05 0.0 24.85 4.57
0.1 0.0 18.84 4.33
0.15 0.0 15.33 4.14
0.2 0.0 12.83 3.96
(b) AudioSeal-B
$\ell_{\infty}$ Bound FNR SNR ViSQOL
0.05 0.15 24.85 4.55
0.1 0.25 18.83 4.29
0.15 0.45 15.33 4.08
0.2 0.55 12.85 3.91
(c) Timbre
$\ell_{\infty}$ Bound FNR SNR ViSQOL
0.05 0.0 25.68 4.79
0.1 0.07 19.66 4.58
0.15 0.14 16.14 4.39
0.2 0.21 13.66 4.22
(d) WavMark
$\ell_{\infty}$ Bound FNR SNR ViSQOL
0.05 0.0 25.96 4.81
0.1 0.67 19.63 4.70
0.15 1.0 16.18 4.58
0.2 1.0 14.52 4.52
Table 8: Detection results under white-box removal attack on FMA.
(a) AudioSeal
SNR 20 30 40 50 60
FNR 1.00 0.85 0.35 0.00 0.00
(b) AudioSeal-B
SNR 20 30 40 50 60
FNR 1.00 1.00 0.95 0.50 0.00
(c) Timbre
SNR 20 30 40 50 60
FNR 1.00 0.95 0.85 0.45 0.10
(d) WavMark
SNR 20 30 40 50 60
FNR 1.00 1.00 1.00 0.50 0.40
Table 9: Detection results under white-box forgery attack on FMA.
(a) AudioSeal
SNR 20 30 40 50 60
FPR 1.00 1.00 1.00 1.00 1.00
(b) AudioSeal-B
SNR 20 30 40 50 60
FPR 1.00 0.95 0.90 0.40 0.30
(c) Timbre
SNR 20 30 40 50 60
FPR 1.00 1.00 0.90 0.50 0.20
(d) WavMark
SNR 20 30 40 50 60
FPR 1.00 1.00 1.00 1.00 1.00
A.11 More Results on AudioMarkData
We applied I-FGSM as an additional white-box attack in Table 10. In Table 11, we show the results for composed no-box perturbations including EnCodeC with 24kHz, MP3 with 16kbps, and Gaussian noise with SNR of 20dB.
Table 10: Results for I-FGSM on AudioMarkData.
(a) Watermark removal (FNR)
SNR AudioSeal AudioSeal-B Timbre WavMark
20 1.00 1.00 1.00 1.00
30 0.75 1.00 1.00 1.00
40 0.25 0.90 0.85 0.50
50 0.00 0.40 0.45 0.50
60 0.00 0.15 0.05 0.15
(b) Watermark forgery (FPR)
SNR AudioSeal AudioSeal-B Timbre WavMark
20 1.00 1.00 1.00 1.00
30 1.00 1.00 1.00 1.00
40 1.00 0.75 1.00 1.00
50 1.00 0.20 0.90 1.00
60 1.00 0.00 0.50 1.00
Table 11: Results for composed no-box perturbations.
(a) EnCodeC + MP3
Method FNR FPR
AudioSeal 0.99 0.00
AudioSeal-B 1.00 0.00
Timbre 0.99 0.00
WavMark 1.00 0.00
(b) MP3 + EnCodeC
Method FNR FPR
AudioSeal 0.95 0.00
AudioSeal-B 1.00 0.00
Timbre 0.96 0.00
WavMark 1.00 0.00
(c) Gaussian noise + MP3
Method FNR FPR
AudioSeal 0.09 0.00
AudioSeal-B 0.65 0.00
Timbre 0.38 0.00
WavMark 1.00 0.00